OpenStudy (anonymous):

soybean meal is 14% protien cornmeal is 7% protein .how many pounds of each should be mixed together in order to get 280 lbs mixture that is 11% protein?

OpenStudy (anonymous):

.14s+.07c=.11

OpenStudy (anonymous):

lets write a little equation here: we will call the soybean meal: S , which of course means that 280-S=Amount of corn meal Now we have : 14S+.07(280-S)=.11*280 .14S+19.6-.07S=30.8 .07S=30.8-19.6 .07S11.2 S=11.2/.07 Which gives us: S=160 Then : 280-160=120 for the cornmeal

OpenStudy (anonymous):

my second equations would have been s+c=280 but i believe lagrange beat me to it lol